# ColorVideoVDP: A visual difference predictor for image, video and display distortions

### v1

@@ 5.2 Ablations
Figure 18: Ablation studies on ColorVideoVDP . The top row shows the base model. Refer to the text for details.
5.3 Synthetic tests
Validation datasets and ablations may not capture the edge cases for which a metric may perform differently than expected. To examine these, we created 14 sets of synthetic test and reference pairs, containing contrast, masking, flicker patterns, and typical distortions. One such example, shown in Fig. 19, demonstrates the metric’s ability to estimate the magnitude of supra-threshold contrast correctly. The lines shown in the plots connect contrast across three color directions that match in perceived magnitude, according to the data of Switkes and Crognale [1999]. The figure shows that while ColorVideoVDP  can correctly predict suprathreshold contrast, HDR-FLIP Andersson et al. [2021] overpredicts the contrast in the red-green and yellow-violet directions. The extensive report for all 14 sets and multiple metrics can be found in the supplementary HTML report.
Figure 19: Testing of supra-threshold contrast matching across color directions. The lines in the plots represent metric responses for the contrast modulated along one of the cardinal directions of the DKL color space (examples shown on the left). The magnitude of the contrast was selected to match the data of Switkes and Crognale [1999]. If a metric correctly predicts the magnitude of chromatic and achromatic contrast, the plots should form horizontal lines. The values for FLIP have been shown as 1-(FLIP mean difference) to make them comparable with other metrics. Note that the perceived magnitudes depend on the spatial frequency and, therefore, the viewing distance.
6 Applications
@@
6.1 Chroma subsampling
@@
Figure 20: The comparison of quality loss due to subsampling of RGB channels (top row), chroma channels of Yxy colorspace (middle row), and chroma channels of YCbCr color space (bottom row). The plots show the predictions of ColorVideoVDP  (1st column) and SSIM (2nd column), the latter computed either on luma or RGB channels. SSIM fails to predict substantial quality loss when the chroma channels are severely subsampled.
6.2 Setting display color tolerance specifications
@@
Figure 21: (a) Spectral power distributions for baseline (red solid line) and modified curves (green dashed line), (b) Proof image shown on display with baseline primaries, (c) Proof image shown on display with modified primaries. © caminandes.com
@@
Figure 22: (a) JOD scores vs spectral peak shift for R, G, B primaries. (b) JOD scores vs. FWHM change from baseline for R, G, B primaries. Error bars represent maximum and minimum JOD scores among 4 scenes for each case.
6.3 Observer metamerism and variability
Observer metamerism has been an issue for wide-color-gamut displays. As wide color gamut displays typically have spectrally narrower primaries, it is increasingly likely that, when calibrated for a standard observer, they will appear color-inaccurate for individuals who deviate from this profile Bodner et al. [2018]; Hung [2019]. The severity of this observer metamerism depends on both the content being shown and the display characteristics, and can be evaluated by using ColorVideoVDP .
To demonstrate this, we calculated JOD scores for the ‘Wiki’ scene, simulating three different display primary spectral profiles and 11 different observer functions. The evaluated displays were Sony BVM32 CRT (100% of sRGB), Eizo ColorEdge 3145 (99% P3), and laser primaries (100% ITU-R BT.2020 color gamuts). Laser spectra were synthetically generated. Spectral power distributions for all primaries are shown in Fig. 23 (top). To simulate individual differences in color vision, we used Asano and Fairchild’s 10 categorical observers Asano and Fairchild [2020], as well as the 1931 ${}^{\circ}$ standard observer used for reference. These 10-observer functions were derived as representative means for a color-normal population based on an individual colorimetric observer model Asano et al. [2016]. The results are shown in Fig. 23 (bottom). As expected, the narrower spectra incur a larger metameric error for non-standard observers, which is reflected in the higher $\Delta$JOD values. ColorVideoVDP  can be used to estimate the risk of metameric error for populations on content, and to evaluate or optimize solutions for this problem.
Figure 23: (Top) Spectral power distributions for three types of displays used to simulate observer metamerism (see Sec. 6.3). (Bottom) JOD scores for all observers relative to 1931 ${}^{\circ}$ standard observer for the ‘Wiki’ scene, for three display types.
7 Conclusions
@@
An important aspect of our metric is its extensive calibration. Our metric is simultaneously calibrated on 3 large datasets containing a variety of distortion types, which included our novel XR-DAVID psychophysical video quality dataset, containing a range of display hardware-oriented artifacts critically important to the development of future displays. We demonstrated that our metric generalizes well to unseen datasets.
@@
Finally, ColorVideoVDP  is efficiently implemented to run on a GPU. This aspect is increasingly important in the modern display landscape, as resolution, frame rate, field of view, and bit-depth continue to increase, adding complexity to analyzed content.
Acknowledgements
@@
The authors would like to thank Minjung Kim, Trisha Lian, and Romain Bachy for the insightful discussions and Xin Li and Curtis Torrey for help with the hardware setup. We also thank the research assistant team: Cameron Wood, Sydnie Gregory, Ka Yan Wat, Julie Stevens, and Helen Ayele for help running the user study and Sara Kenley, Jessica Grey, and Eloise Moore for assistance with logistics.
Supplementary for ColorVideoVDP: A visual difference predictor for image, video and display distortions
This supplementary document describes:
  • the procedure used to convert from the cone contrast units used in castleCSF to the DKL contrast used in ColorVideoVDP  (Sec. S1);
  • the contrast encoding (Sec. S2) and contrast masking functions (Sec. S3) that were considered for ColorVideoVDP  and are included in the ablations;
  • several mathematical techniques that were necessary to make the model differentiable (Sec. S4);
  • timings of ColorVideoVDP  compared to VMAF (Sec. S5);
  • the experiment used to obtain a JOD quality scaling for the LIVEHDR dataset (Sec. S6);
  • details on the content used for XR-DAVID dataset (Sec. S7).
S1 Contrast sensitivity in the DKL color space
The castleCSF Anonymous [2024] model predicts sensitivity in different contrast units than those used by ColorVideoVDP . To use castleCSF in ColorVideoVDP , we need to convert between the two contrast units. The contrast conversion procedure is similar to the one used in Kim et al. [2021].
The sensitivity used in castleCSF is defined as the inverse of cone contrast:
$S_{\text{cc}}=\sqrt{3}\left(\left(\frac{\Delta\mathfrak{L}}{\mathfrak{L}_{0}}\right)^{2}+\left(\frac{\Delta\mathfrak{M}}{\mathfrak{M}_{0}}\right)^{2}+\left(\frac{\Delta\mathfrak{S}}{\mathfrak{S}_{0}}\right)^{2}\right)^{-0.5}\,,$ (S1)
where $\Delta\mathfrak{L}$, $\Delta\mathfrak{M}$, and $\Delta\mathfrak{S}$ are the amplitudes of the cone responses for the stimuli and $\mathfrak{L}_{0}$, $\mathfrak{M}_{0}$, and $\mathfrak{S}_{0}$ are the cone responses for the corresponding background. ColorVideoVDP  encodes contrast for the three cardinal dimensions of the DKL space ($c\in\{\text{Ach},\text{RG},\text{YV}\}$) as:
$C_{c}=\frac{\mathcal{L}_{c}}{L}\,,$ (S2)
[…]
$S_{\text{DKL}{,c}}=C^{-1}_{c}=\frac{L}{\mathcal{L}_{c}}\,,$ (S3)
To convert $S_{\text{cc}}$ into $S_{\text{DKL}{,c}}$, we first obtain from castleCSF the sensitivity $S_{\text{cc}}$ along each cardinal color direction ($c$) of the DKL color space. Then, we find the contrast $C_{c}$ that results in the threshold cone contrast corresponding to the predicted sensitivity ($S^{-1}_{\text{cc}}$ since the threshold contrast is the inverse of the sensitivity). As there is no closed-form solution, $C_{c}$ needs to be found by non-linear root finding. In the optimization loop, we allow the DKL contrast $C_{c}$ to change only along the given cardinal color direction. This calculation is repeated for each spatial frequency, luminance, and color direction. The sensitivities in the DKL contrast space are precomputed and stored in a look-up table for later use by ColorVideoVDP , as explained in Sec. 3.4 of the main paper.
S2 Contrast encoding
Multiplicative contrast normalization
Once we have separated the band-limited contrast in each frequency band and visual channel (two achromatic and two chromatic), we want to encode it in a way that correlates well with the perceived magnitude of the contrast. To this end, many visual models employ the normalization by the contrast sensitivity function $S(\cdot)$
$C^{\prime}=C\,S(\rho_{b},L_{\text{bkg}},c)\,,$ (S4)
where $C$ is the physical and $C^{\prime}$ is encoded contrast, $\rho_{b}$ is the peak frequency of the band $b$, $c$ is the channel index, and $L_{\text{bkg}}$ is the background luminance. Since contrast can be expressed as $C=\nicefrac{{\Delta{L}}}{{L_{\text{bkg}}}}$, we can write:
$C^{\prime}=C\,S(\rho_{b},L_{\text{bkg}},c)=\frac{\Delta{L}}{L_{\text{bkg}}}\frac{L_{\text{bkg}}}{\Delta{L_{\text{thr}}}}=\frac{\Delta{L}}{\Delta{L_{\text{thr}}}}\,,$ (S5)
where $\Delta{L_{\text{thr}}}$ is the luminance difference corresponding to the detection threshold. It follows that $C^{\prime}$ encodes multiples of the detection threshold. Notably, it is equal to 1 when the contrast $C$ is exactly at the threshold value. This property is necessary for any contrast encoding employed by our method because masking models, discussed in the next section, rely on it. The encoding is plotted in the left panel of Fig. S1.
The above contrast normalization by sensitivity brings several benefits. Daly Daly [1993] showed that such a normalization is necessary to unify masking predictions across spatial frequencies. Peli et al. Peli et al. [1991] demonstrated that this normalization accounts for contrast matching across luminance levels.
Additive contrast normalization
[…]
$C^{\prime}_{K}=C-\frac{1}{S(\rho_{b},L_{\text{bkg}},c)}\,.$ (S6)
Note that the inverse of the contrast sensitivity in this equation is the threshold visibility contrast. Kulikowski Kulikowski [1976] demonstrated that perceived contrast is matched across luminance levels when the corresponding $C^{\prime}_{K}$ is matching (the sensitivity $S$ in the equation varies across luminance levels). Kulikowski’s model assumes that the detection thresholds are caused by additive (neural) noise, and therefore, the visual system ”eliminates” the noise by subtracting it from the contrast signal.
To support both contrast polarities, we modify the above formula as follows:
$C^{\prime}=\textrm{sgn}{(C)}\max\left\{g\left(|C|-\frac{1}{S(\rho_{b},L_{\text{bkg}},c)}\right)+1,0\right\}\,,$ (S7)
where $g$ is the gain that controls the strength of the contrast above the threshold. As in the previous case, the constant 1 is used to ensure that the encoded contrast has a value of 1 at the threshold. This is an important modification as it allows us to represent contrasts below the detection threshold. The sign function ensures that we preserve the polarity of the contrast. The encoding is plotted in the right panel of Fig. S1.
Figure S1: Mapping from input physical contrast to encoded contrast for multiplicative (left) and additive (right) encoding.
To check how well each contrast encoding represents perceived contrast, in the following sections we generate contrast-matching predictions and compare them with those reported in the literature.
[…]
S2.1 Matching chromatic and achromatic contrast
[…]
To improve the accuracy of these matches, we introduced corrections into the contrast encoding equation for the multiplicative contrast encoding:
$C^{\prime}=m_{c}\,C\,S(\rho_{b},L_{\text{bkg}},c)\,,$ (S8)
where $m_{c}=\begin{bmatrix}1&1&1.45&0.95\end{bmatrix}$ for the color channels:
$c\in\{\text{Ach}\mathcal{S},\text{Ach}\mathcal{T},\text{RG},\text{YV}\}$ (S9)
corresponding to the achromatic sustained, achromatic transient, chromatic red-green, and chromatic yellow-violet channels. The additive contrast encoding needs to include an additional gain factor:
$C^{\prime}=\textrm{sgn}{(C)}\max\left\{g\,m_{c}\left(|C|-\frac{1}{S(\rho_{b},L_{\text{bkg}},c)}\right)+1,0\right\}\,,$ (S10)
where $m_{c}=\begin{bmatrix}1&1&1.7&0.237\end{bmatrix}$. The adjusted contrast encoding equations are plotted as continuous lines Fig. S2.
S2.2 Matching contrast across frequencies
Figure S3: The lines connecting the matching magnitude of contrast across frequencies, as predicted by the multiplicative (left) and additive (right) contrast encoding.
The seminal paper of Georgeson and Sullivan [1975] demonstrated contrast constancy — the observation that the perceived magnitude of contrast differs across spatial frequencies when the contrast is small (near the detection threshold) but when the contrast is far above the threshold, there is little difference in the perceived magnitude of contrast. This property is better captured by the additive contrast encoding model, shown in the right panel of Fig. S3. This contrast constant effect is easy to observe in everyday life — objects seen from close and far distances will fall into very different spatial frequency ranges. Yet, we do not observe changes in object appearance as we move closer or further away from them. The multiplicative contrast encoding technique does not directly model the property of contrast constancy.
S2.3 Matching contrast across luminance
Multiple authors investigated whether contrast constancy generalizes across luminance levels Kulikowski [1976]; Georgeson and Sullivan [1975]; Hess [1990]; Peli et al. [1991]; Peli [1995] and they all observed quite a significant deviation from contrast constancy, especially when the luminance drops significantly below photopic levels. However, there is no agreement on how to model such a deviation from contrast constancy. Both Kulikowski [1976] and Georgeson [1991] proposed additive contrast encoding to explain their contrast matching data. However, Peli [1995] demonstrated that a multiplicative model better explains the data when contrast is seen naturally rather than using a dichoptic presentation used in other studies (each eye sees a different luminance). The predictions for both models are shown in Fig. S4.
Figure S4: The lines connecting the matching magnitude of contrast across luminance, as predicted by the multiplicative (left) and additive (right) contrast encoding.
S3 Contrast masking
The main purpose of the masking model is to transform physical differences in contrast between two images or frames into perceived differences. Here, consider three models of masking: a model based on contrast transducer functions, such as the one proposed by Watson and Solomon Watson and Solomon [1997], a model based on mutual masking, such as the one proposed in Daly’s original VDP Daly [1993], and the similarity formula, used in SSIM and other metrics.
Contrast transducer
The difference between two band-limited images in Watson and Solomon’s Watson and Solomon [1997] masking model is expressed as a difference between two contrast transducer functions:
$D_{b,c,f}(\bm{x})=\left|t(C_{b,c,f}^{\prime\text{test}}(\bm{x}))-t(C_{b,c,f}^{\prime\text{ref}}(\bm{x}))\right|$ (S11)
where $b$ is the index of the spatial frequency band, $c$ is the channel index (two achromatic and two chromatic), and $f$ is the frame index. $C_{b,c,f}^{\prime\text{test}}$ and $C_{b,c,f}^{\prime\text{ref}}$ correspond to the encoded contrast in the test and reference images (refer to the main paper). The transducer function is formulated as:
$t(C^{\prime}_{b,c,f})=g_{\text{T}}\frac{\textrm{sgn}(C^{\prime}_{b,c,f}(\bm{x}))\left|C^{\prime}_{b,c,f}(\bm{x})\right|^{p}}{0.2+\sum_{i}k_{i,c}\left(\left|C^{\prime}_{b,i,f}\right|^{q_{c}}*g_{\sigma_{\text{sp}}}\right)(\bm{x})}\,,$ (S12)
where $p$ and $q_{c}$ are the parameters of the model. The masking parameter $q_{c}$ is set separately for each channel (two achromatic and two chromatic channels). $g_{\text{T}}$ is the gain of the transducer that lets us control the range of visual difference values. The constant of 0.2 was selected so that the facilitation (the dip in the contrast discrimination function) coincides with the measurements, as explained in Sec. S3.2. The expression in the denominator pools contrast in a local spatial neighborhood by convolving with a Gaussian kernel $g_{\sigma_{\text{sp}}}$ with the standard deviation of $\sigma_{\text{sp}}$. The sum in this expression pools contrast across channels according to the cross-masking coefficient $k_{c,i}$. $C^{\prime}$ is the encoded contrast, as explained in Sec. S2.
Mutual masking
Alternatively, the difference between two band-limited images can be expressed using the mutual masking model, as proposed by Daly Daly [1993]:
$D_{b,c,f}(\bm{x})=\frac{\left|C^{\prime\,\textrm{test}}_{b,c,f}(\bm{x})-C^{\prime\,\textrm{ref}}_{b,c,f}(\bm{x})\right|^{p}}{1+(C^{\textrm{mask}}_{b,c,f}(\bm{x}))^{q_{c}}}$ (S13)
First, the mutual masking of test and reference bands is calculated as [Daly, 1993, p.192]:
$C^{\text{mm}}_{b,c,f}(\bm{x})=\min\left\{\left|C^{\prime\,\textrm{test}}_{b,c,f}(\bm{x})\right|,\left|C^{\prime,\textrm{ref}}_{b,c,f}(\bm{x})\right|\right\}\,.$ (S14)
Then, similarly to the previous model, the mutual masking signal is pooled in a small local neighborhood by convolving with a Gaussian kernel $g_{\sigma_{\text{sp}}}$, and combined across channels, accounting for cross-channel masking:
$C^{\textrm{mask}}_{b,c,f}(\bm{x})=\sum_{i}k_{i,c}(C^{\text{mm}}_{b,i,f}*g_{\sigma_{\text{sp}}})(\bm{x})\,.$ (S15)
One shortcoming of the mutual masking model is that it does not account for self-masking — the case in which the reference contrast is 0 and the test contrast is attenuated when its value is high. In practice, $D_{b,c,f}$ can reach very high values (over 10 000) when the sensitivity is high. This is unrealistic behavior as neurons cannot encode values of such high dynamic ranges. For that reason, we need to limit the range of contrast difference values with a smooth clamping function:
$\hat{D}_{B,c,f}(\bm{x})=\frac{D_{\text{max}}\,D_{B,c,f}(\bm{x})}{D_{\text{max}}+D_{B,c,f}(\bm{x})}\,,$ (S16)
where $D_{\text{max}}$ is the maximum value that the visual difference can attain.
Similarity
The alternative masking model, found in SSIM and many other metrics, can be formulated as:
$D_{b,c,f}(\bm{x})=D_{\text{max}}-D_{\text{max}}\frac{2\,\left|C^{\prime\,\textrm{test}}_{b,c,f}(\bm{x})\right|\,\left|C^{\prime\,\textrm{ref}}_{b,c,f}(\bm{x})\right|+\epsilon}{\left(\hat{C}^{\textrm{test}}_{b,c,f}\right)^{2}(\bm{x})+\left(\hat{C}^{\textrm{ref}}_{b,c,f}\right)^{2}(\bm{x})+\epsilon}\,,$ (S17)
where $\epsilon=D_{\text{max}}-1$ is selected so that the resulting visual difference is 1 when the test contrast is at the detection threshold ($C^{\prime\,\textrm{test}}_{b,c,f}=1$) and the reference contrast is 0. $D_{\text{max}}$ controls the maximum value of the visual difference. The standard formula typically uses the same values in both nominator and denominator. Here, we modify the denominator so that it contains masking signal $\hat{C}^{\textrm{test}}_{b,c,f}$ and $\hat{C}^{\textrm{ref}}_{b,c,f}$ associated with the test and reference images. For the test image
$\hat{C}^{\textrm{test}}_{b,c,f}(\bm{x})=\sum_{i}k_{i,c}(C^{\prime\textrm{test}}_{b,i,f}*g_{\sigma_{\text{sp}}})(\bm{x})\,.$ (S18)
and the masking signal for the reference frame is computed analogously.
Figure S5: Model response for self-masking — when the reference contrast is 0 (is a uniform field) but the test contrast (x-axis) and sensitivity (colors) vary. The predictions are shown for all combinations of contrast coding (columns) and masking models (rows). The dashed horizontal line at $D=1$ represents the difference at the detection threshold.
S3.1 Self-masking
To give more insights into the three masking models explained above and the two contrast encodings, we plot the model predictions for the case of self-masking in Fig. S5. Self-masking is the case in which the contrast (in the test) is masked by itself and is not influenced by the contrast in the reference image (the reference contrast is). This scenario may appear e.g. when the reference image is a uniform field, and the test image contains a pattern that we want to detect. All the plots in Fig. S5 show that the visual difference is equal to 1 (the dashed horizontal line) when the test contrast is at the detection threshold (equal to $\nicefrac{{1}}{{S}}$). The additive contrast encoding (right column) will result in small contrast being ignored when the sensitivity is low, while the multiplicative encoding will compress contrast below the detection threshold. The mutual masking and similarity models are remarkably similar to each other in terms of self-masking.
Figure S6: Model response for contrast masking. The reference image contains a certain physical contrast $C^{\text{mask}}$ (x-axis), and the test image contains the physical contrast of $C^{\text{mask}}+{\Delta}C^{\text{test}}$ (y-axis). The contour lines denote the response of the model (blue for the smallest and yellow for the largest values). The red circles denote the contrast masking measurements from [Foley, 1994, Figure 3b, data for KMF]. The axes roughly correspond to those found in the example in Fig. S7.
Figure S7: The example of typical stimuli used in contrast masking experiments. A test Gabor patch is superimposed (added) on the background of a sinusoidal grating of the same frequency (a masker). As the contrast of the masker is increased (towards the right), a higher test contrast of the Gabor patch is needed to detect it.
S3.2 Contrast masking
The standard contrast masking experiment involves showing a test pattern superimposed on top of a masking pattern, such as the one shown in Fig. S7. We simulate the same scenario for our six combinations of masking models and contrast encodings and show the results in Fig. S6. In each plot, we include the measurements of contrast masking from Foley [1994] as red dots. We expect the contour lines to follow the curve formed by those measurements. In the plots, we can see that only the transducer models the facilitation that makes patterns easier to detect when the masker is near the detection threshold. This is shown as a small dip in the measurements by Foley [1994]. Mutual masking and similarity models show similar behavior.
S4 Training considerations
While the contrast encoding and masking models proposed in the previous section are well-defined for an arbitrary contrast, these models are not differentiable out-of-the-box and, therefore, cannot be used in gradient-based optimization. To make them differentiable, we replace the sign function with a hyperbolic tangent function:
$\textrm{sgn}(C)\approx\tanh(10000\,C)\,.$ (S19)
The exponential function is not differentiable at 0, therefore, we need to approximate it as:
$|C|^{p}\approx(|C|+\epsilon)^{p}-\epsilon^{p}\,,$ (S20)
where $\epsilon=0.00001$.
S5 Quality metric timings
We measured the times required to process 50 video frames at resolutions ranging from 720p to 4K. The measurements were averaged across 5 runs and excluded the times required to load and decode the frames. The measured times, illustrated in Fig. S8, show that ColorVideoVDP  processes large videos faster than VMAF (note however, that VMAF was running on a CPU) and in about the same time as FovVideoVDP. It must be noted, however, that unlike ColorVideoVDP both VMAF and FovVideoVDP operate only on luma or luminance and ignore color channels. VMAF features are much less expensive to extract and FovVideoVDP processes only 2 visual channels (sustained and transient achromatic) compared to the 4 channels processed by ColorVideoVDP .
Figure S8: The processing times for 50-frame video of different resolutions (x-axis). ColorVideoVDP  and FovVideoVDP were run on an Nvidia Quadro RTX 8000 GPU, while VMAF was run on 4 cores of an Intel Xeon Gold 5218 CPU @ 2.30 GHz. The dashed vertical lines indicate the standard video resolutions, from 720p to 2160p (4K). All the reported times were averaged across five runs.
S6 LIVE HDR dataset alignment experiment
In order to enable training on multiple datasets, it is essential to bring their quality scores to a common scale. This ensures that quality scores from one dataset are directly comparable and equivalent to those from another dataset. While both XR-DAVID and UPIQ employ the same quality units (JODs), LIVE HDR was originally collected using the mean-opinion-score (MOS) units. To account for this discrepancy, we ran an additional experiment that let us scale LIVE HDR MOS values in the JOD units.
We first calculated differences of mean opinion scores (DMOS) by subtracting the score of the reference (highest quality video) from each condition. Next, in order to maximize the sampling of quality across the dataset, we carefully selected 10 videos that contained a wide distribution of MOS values. For each of these videos, we then selected 6 different levels of distortion. We used this content to conduct an experiment using the same experimental procedure as for XR-DAVID, as explained in the main paper. Display and environment settings were adapted to mimic those used in the original LIVE HDR study. 12 participants took part in the study, and completed 10 repetitions (batched) of ASAP sampling, for a total of $10\times 6\times 10=600$ trials each. The resulting JOD values for each condition were used to find a linear mapping from the LIVE HDR DMOS scores to our experimentally derived JOD scores. The resulting linear fit, illustrated in Fig. S9, demonstrates a satisfactory level of accuracy. Thus, we applied this mapping to the entire LIVE HDR dataset, effectively rescaling it to JOD units.
Figure S9: We conducted a subjective study, obtaining JOD scores for a subset of videos from the LIVE HDR dataset. A linear fit was obtained as follows: $\textnormal{JOD}=-0.054*\textnormal{DMOS}+0.037$, which was deemed acceptable. JOD values are shifted from 0 to 10 by convention.
S7 XR-DAVID content
Thumbnails of select scenes used in the XR-DAVID dataset are shown in Fig. S10. The source data for each video is shown in Table S1. Note that for longer videos, sections of approximately $5$ seconds from the start of each video were used in the study, with the exception of Caminandes where a shot from approximately 1:54-2:00 of ”Caminandes 3: Llamigos” was used. Videos were downsampled to a resolution of $910\times 540$ using bilinear interpolation, which produces an effective visual resolution of $\approx 40$ppd when seen by users from a distance of $28.9$’. Prior to display, videos were upscaled using nearest-neighbor interpolation by a factor of $2$ to $1920\times 1080$ to match the native resolution of the display. Detailed numerical results of our experiment are shown in Fig. S11 for each of the $3$ levels of intensity per artifact studied.
Table S1: XR-DAVID assets
Video Source Duration
Bonfire Kindel Media [link] 5.6s
Business Kindel Media [link] 5.6s
Caminandes Blender Foundation [link] 4.8s
Couple RDNE Stock project [link] 5.6s
Dance Anna Shvets [link] 5.6s
Emojis emirkhan bal [link] 5.6s
Foliage German Korb [link] 5.6s
Icons Generated by the authors 5.6s
Panel Generated by the authors 5.6s
Cellphone Lina Fresco [link] 5.6s
River Theresa. Nguyen [link] 3.7s
Snow SwissHumanity Stories [link] 5.6s
River Theresa. Nguyen [link] 3.7s
VR Generated by the authors 5.6s
Wiki Generated by the authors 5.2s
Figure S10: Thumbnails of 8 out of 14 scenes used in the XR-DAVID study. The bottom-right plot shows the color gamut coverage of all the pixels in the reference videos on a CIELAB a* b* plane. The thumbnails of the videos containing human subjects could not be included due to institutional policy.
Figure S11: The results of our XR-DAVID color video quality experiment. For each base video, 8 different artifacts were studied at 3 intensity levels (1:top, 2:middle, and 3:bottom). The responses are scaled on a single perceptual JOD scale, counting down from 10 by convention. Increasing magnitudes of perceived distortion can be observed at stronger distortion levels (top-to-bottom). In addition, large differences in artifact visibility can be observed across content (columns).

### v2

@@ 5.2. Ablations
Figure 18. Ablation studies on ColorVideoVDP. The top row shows the base model. Refer to the text for details.
5.3. Synthetic tests
Validation datasets and ablations may not capture the edge cases for which a metric may perform differently than expected. To examine these, we created 14 sets of synthetic test and reference pairs, containing contrast, masking, flicker patterns, and typical distortions. One such example, shown in Fig. 19, demonstrates the metric’s ability to estimate the magnitude of supra-threshold contrast correctly. The lines shown in the plots connect contrast across three color directions that match in perceived magnitude, according to the data of Switkes and Crognale (1999). The figure shows that while ColorVideoVDP can correctly predict suprathreshold contrast, HDR-FLIP (Andersson et al., 2021) overpredicts the contrast in the red-green and yellow-violet directions. The extensive report for all 14 sets and multiple metrics can be found in the supplementary HTML report.
Figure 19. Testing of supra-threshold contrast matching across color directions. The lines in the plots represent metric responses for the contrast modulated along one of the cardinal directions of the DKL color space (examples shown on the left). The magnitude of the contrast was selected to match the data of Switkes and Crognale (1999). If a metric correctly predicts the magnitude of chromatic and achromatic contrast, the plots should form horizontal lines. The values for FLIP have been shown as 1-(FLIP mean difference) to make them comparable with other metrics. Note that the perceived magnitudes depend on the spatial frequency and, therefore, the viewing distance.
6. Applications
@@
Figure 20. The comparison of quality loss due to subsampling of RGB channels (top row), chroma channels of Yxy colorspace (middle row), and chroma channels of YCbCr color space (bottom row). The plots show the predictions of ColorVideoVDP (1st column) and SSIM (2nd column), the latter computed either on luma or RGB channels. SSIM fails to predict substantial quality loss when the chroma channels are severely subsampled.
6.1. Chroma subsampling
@@
6.2. Optimization
ColorVideoVDP is fully differentiable, and it can be potentially used as an image loss function when training deep neural networks. However, a good quality metric may not serve as a good loss function (Mustafa et al., 2022), in particular when the features used by a metric are not surjective (Ding et al., 2021). Moreover, complex functions can make the landscape of the loss highly non-convex and impede convergence. For that reason, a popular ”perceptual loss” (Johnson et al., 2016) must be combined with L1 or L2 to ensure convergence.
We found that when ColorVideoVDP is used as an image loss function (without L1 or L2 term) in an image reconstruction task (see (Ding et al., 2021, Sec. 3)), it converges well when the reconstructed image is close to the reference. However, the convergence is worse when the test image is affected by random noise or is far from the reference. This problem is alleviated when ColorVideoVDP is used when optimizing low-dimensional problems, such as the selection of encoding parameters in video streaming. In the next section, we show an example of ColorVideoVDP used as a loss function in a high-dimensional problem.
6.3. Adaptive chroma subsampling
Figure 21. Adaptive chroma subsampling removes high frequencies from chromatic Cb and Cr planes (2nd and 3rd column) of the YCbCr color space without affecting the appearance of the color image (1st column). Such adaptive subsampling can effectively improve image compression without the loss of chromatic details.
Traditional chroma subsampling, as demonstrated in Sec. 6.1, globally reduces the resolution (sampling rate) of the chromatic planes regardless of image content. However, the visibility of chromatic distortion will strongly depend on image content, and therefore, we argue that chroma subsampling should be adaptive and vary across an image. Here, we show that we can use ColorVideoVDP as a differential loss to adaptively remove high frequencies from the image chroma channels while accounting for all factors that contribute to the visibility of the introduced changes. To do this, we optimize for the color image $I$:
(19) $\operatorname*{argmin}_{I}||I-I_{\text{org}}||_{\text{cvvdp}}+\lambda\left||\nabla I_{\text{Cb}}\right||_{1}+\lambda||\nabla I_{\text{Cr}}||_{1}$
where $I_{\text{org}}$ is the original image, $\nabla I_{\text{Cb}}$ and $\nabla I_{\text{Cr}}$ are the gradients of the Cb and Cr planes of the YCbCr color space and $\lambda$ is a regularization constant. $||\cdot||_{\text{cvvdp}}$ denotes ColorVideoVDP used as a differentiable loss and $||\cdot||_{1}$ is an L1 norm. Fig. 21 shows an example of such adaptive chroma subsampling. In this particular example, we could reduce the size of a PNG file by 10% with no impact on image quality.
Figure 22. (a) Spectral power distributions for baseline (red solid line) and modified curves (green dashed line), (b) Proof image shown on display with baseline primaries, (c) Proof image shown on display with modified primaries. © caminandes.com
6.4. Setting display color tolerance specifications
@@
Figure 23. (a) JOD scores vs spectral peak shift for R, G, B primaries. (b) JOD scores vs. FWHM change from baseline for R, G, B primaries. Error bars show maximum and minimum JOD scores among 4 scenes for each case.
6.5. Observer metamerism and variability
Observer metamerism has been an issue for wide-color-gamut displays. As wide color gamut displays typically have spectrally narrower primaries, it is increasingly likely that, when calibrated for a standard observer, they will appear color-inaccurate for individuals who deviate from this profile (Bodner et al., 2018; Hung, 2019). The severity of this observer metamerism depends on both the content being shown and the display characteristics, and can be evaluated by using ColorVideoVDP.
To demonstrate this, we calculated JOD scores for the ‘Wiki’ scene, simulating three different display primary spectral profiles and 11 different observer functions. The evaluated displays were Sony BVM32 CRT (100% of sRGB), Eizo ColorEdge 3145 (99% P3), and laser primaries (100% ITU-R BT.2020 color gamuts). Laser spectra were synthetically generated. Spectral power distributions for all primaries are shown in Fig. 24 (top). To simulate individual differences in color vision, we used Asano and Fairchild’s 10 categorical observers (2020), as well as the 1931 2 standard observer used for reference. These 10-observer functions were derived as representative means for a color-normal population based on an individual colorimetric observer model (Asano et al., 2016). The results are shown in Fig. 24 (bottom). As expected, the narrower spectra incur a larger metameric error for non-standard observers, which is reflected in the higher $\Delta$JOD values. ColorVideoVDP can be used to estimate the risk of metameric error for populations on content, and to evaluate or optimize solutions for this problem.
Figure 24. (Top) Spectral power distributions for three types of displays used to simulate observer metamerism (see Sec. 6.5). (Bottom) JOD scores for all observers relative to 1931 2 standard observer for the ‘Wiki’ scene, for three display types.
7. Conclusions
@@
An important aspect of our metric is its extensive calibration. Our metric is simultaneously calibrated on 3 large datasets containing a variety of distortion types, which included our novel XR-DAVID video quality dataset, containing a range of display hardware-oriented artifacts critically important to the development of future displays. We demonstrated that our metric generalizes well to unseen datasets.
@@
Finally, ColorVideoVDP is efficiently implemented to run on a GPU. This aspect is increasingly important in the modern display landscape, as resolution, frame rate, field of view, and bit-depth continue to increase, adding complexity to analyzed content. It is also fully differentiable, making it possible to use it as a loss function in optimization problems.
Acknowledgements.
@@
We thank Dounia Hammou for her insightful comments. We thank Cameron Wood, Eloise Moore, Helen Ayele, Sydnie Gregory and Ka Yan Wat for help running the user study. Thanks go to Curtis Torrey and Xin Li for assistance with experimental hardware, and to Jacqueline Beyrouty for help with logistics.